# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v1

@@ 14 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$1 0 -3 0 -1
Appendix B Experiments: More Details and Further Results
B.1 Datasets
@@
Following [Tzeng et al., 2020], we consider the following widely studied real-world signed networks. WoW-EP8 (W8) [Kristof et al., 2020] represents interactions among authors in the 8th EU Parliament legislature, where edge signs indicate collaboration or competition. Bitcoin (BTC) [Leskovec and Krevl, 2014] is a trust-distrust network of users trading on the Bitcoin OTC platform. WikiVot (WikiV) [Leskovec and Krevl, 2014] records positive and negative votes for Wikipedia admin elections. Referendum (REF) [Lai et al., 2018] captures tweets about the 2016 Italian constitutional referendum, with edge signs indicating whether users share the same stance. Slashdot (SD) [Leskovec and Krevl, 2014] is a friend-foe network from the Slashdot Zoo feature. WikiCon (WikiC) [Kunegis, 2013] tracks positive and negative interactions between users editing English Wikipedia. Epinions (EP) [Leskovec and Krevl, 2014] represents the trust-distrust relationships in the Epinions online social network. WikiPol (WikiP) [Maniu et al., 2011] captures interactions among users editing Wikipedia pages on political topics.
B.2 Baselines
@@
For SCG, we use the public implementation from [Ruo-Chun Tzeng, 2020]. For KOCG, we use the public implementation from [Lingyang Chu, 2016] with default hyperparameters: $\alpha=1/(k-1)$, $\beta=50$ (note that the purpose of this $\beta$ differs from the one used in our paper), and $\ell=5000$. For the spectral methods SPONGE and BNC, we use the public implementations from [Peter Davies and Aldo Glielmo, 2019]. Following [Tzeng et al., 2020], for SPONGE, we evaluate both the unnormalized and symmetric normalized versions and report results for the best-performing method.
B.3 Complete Results on All Datasets
Tables 4-5 report results for polarity and BA-Polarity with $k=2$, $k=4$, and $k=6$. The conclusions for $k=2$ are consistent with those for $k=4$ and $k=6$, discussed in the main paper. According to these results, our method consistently achieves the highest scores or performs competitively with the best scores.
@@ B.3 Complete Results on All Datasets
Tables 6-13 present detailed analyses for all datasets across 12 different aspects, including polarity and BA-polarity. They provide insights into the differences between the clustering solutions computed by each method. In the context of PCD, evaluating clustering quality is inherently subjective, as different aspects capture different aspects of the solution. Generally, there is a trade-off among these aspects, and the objective is to achieve a good balance between them. Below, we define the aspects used in our analysis. We have defined these aspects such that a larger number is better (apart from runtime).
@@
  • BAL = $S_{m}/S_{l}$: The balance factor, where $S_{m}$ and $S_{l}$ are the largest and smallest non-neutral clusters, respectively. Its range it $[0,1]$, with $1$ indicating perfect balance.
@@
  • POL: Polarity, as defined in the main paper.
  • BA-POL: Balance-aware polarity, as defined in the main paper.
  • RT: Runtime of the corresponding method in seconds.
From Tables 6-13, we derive several key insights. Firstly, our method exhibits low standard deviation, indicating robustness to the initial random solution. It consistently finds high-polarity solutions while maintaining better balance than its main competitor, SCG. Unlike some baselines such as KOCG and SPONGE, our method does not enforce excessive balance, ensuring solutions remain both high in polarity and reasonably balanced.

### v2

@@ Appendix C Impact of $\alpha$ and $\beta$
From Prop. 4, we understand the extreme cases of $\beta$: (a) a small negative $\beta$ results in a maximally imbalanced non-neutral clustering (i.e., all objects in one non-neutral cluster), while (b) a large positive $\beta$ makes all objects neutral. For intermediate $\beta\in[\xi_{1},\xi_{2}]$, we analyze the gradient in Eq. 8. Increasing $\beta$ strictly reduces the contribution of object $i$ to each cluster $m\in[k]$, but since the term $-2\beta|S_{m}|$ scales with cluster size, larger clusters become less favorable, promoting balance. If $\beta$ is large enough, it forces $G_{i,m}<0$ for all $m\in[k]$, making neutrality optimal for object $i$. Note that this is more likely for low-degree objects, implying that high-degree objects (with clear cluster assignment) are more likely to remain non-neutral, resulting in dense non-neutral clusters. Consequently, increasing $\beta$ leads to smaller (i.e., more neutral objects) and denser non-neutral clusters, while maintaining balanced, as desired.
The parameter $\alpha$ has been studied in prior work [12, 44]. From Eq. 3, $\alpha$ balances maximizing intra-similarities and minimizing inter-similarities, which translates to a trade-off between cohesion within clusters and separation between them. A heuristic choice of $\alpha=1/(k-1)$ was proposed in [44], based on the observation that the number of intra-similarities scale linearly with $k$, while the number of inter-similarities grow quadratically. This choice prevents inter-similarities from dominating the objective. Finally, the term $-\alpha\phi_{i}$ indicates that $\alpha$ influences whether object $i$ becomes neutral, underscoring the need to account for inter-similarities in the objective (as suggested in Section 2.2).
Appendix D Limitations of Polarity
[…]
Appendix E Motivation for Discovering Balanced Communities
From an optimization perspective, the graph-clustering literature recognizes that objectives that simply maximize intra-cluster similarity or minimize inter-cluster similarity often degenerate into trivial solutions in which all objects are assigned to a single (or very few) clusters. A classic instance is the minimum cut, which minimizes inter-cluster similarity yet exhibits this pathology on both signed and unsigned graphs. The usual remedy is to normalize the objective by a quantity that reflects cluster size or degree, giving rise to alternative measures such as the (signed) ratio cut and the normalized cut [11, 10]. Crucially, such normalization can yield solutions whose value with respect to the unnormalized criterion is worse, while better coinciding with the true underlying clusters. In this work we tackle an analogous limitation in the context of PCD, where neutral objects are permitted.
We next motivate—through a few illustrative examples—why clusterings that are more balanced in size are frequently better aligned with ground-truth clusterings observed in real-world settings.
As discussed in [21], many social environments—from online forums and political systems to scientific institutions—can benefit from balanced community structures in signed networks. Such balance helps ensure that diverse viewpoints or specializations are sufficiently represented and reduces the chance that one perspective overwhelmingly dominates. This, in turn, promotes constructive debate, encourages critical thinking, and helps broaden individuals’ perspectives—ultimately limiting echo chambers and mitigating the spread of misinformation. Below, we provide a few concrete examples.
In market research and product development, balanced communities can provide deeper insights into consumer preferences. By identifying groups of comparable size (rather than one massive consumer segment overshadowing niche but meaningful ones), organizations can more effectively tailor products or marketing strategies to each community, thereby predicting market trends more accurately.
Academic Research Networks can also profit from balanced structures. Imagine a signed network in a university, where positive edges link researchers working on similar topics (e.g., physics, computer science, mathematics), and negative edges indicate differing research domains. Because universities aim to cover a broad range of subjects—each with enough faculty to maintain healthy research output—a balanced partition of the signed network (i.e., each discipline having a solid, non-negligible presence) offers a better representation of the university’s diverse pursuits. In such a setting, neutral nodes might be faculty primarily engaged in administrative roles or other staff members not actively involved in research.
Another example is social networks where communities are formed around specific interests or activities (e.g., online book clubs, gaming groups). These communities grow naturally to a size where interaction remains high, and members can still engage meaningfully with each other, without the network becoming too large or fragmented, i.e., the size of the group grows until it reaches a point where communication remains efficient and personal connections are preserved. As the group grows beyond a certain size, it may split into smaller subgroups, keeping the communities balanced in size.
In all these scenarios, balanced community detection ensures that no single faction’s needs or views go unnoticed. This inclusivity promotes more representative outcomes, strengthens consensus-building, and leads to more robust decisions or insights.
Appendix F Comparison with $\gamma$-polarity Objective from [21]
The $\gamma$-polarity objective introduced in [21] is defined specifically for the case of $k=2$ clusters. In this setting, we have only two non-neutral groups: $S_{1}$ and $S_{2}$. Let $s_{\text{max}}=\max(|S_{1}|,|S_{2}|)$ and $s_{\text{min}}=\min(|S_{1}|,|S_{2}|)$. The $\gamma$-polarity is then defined as
$\frac{N^{+}_{\text{intra}}-N^{-}_{\text{intra}}+N^{-}_{\text{inter}}-N^{+}_{\text{inter}}}{(s_{\text{max}}-s_{\text{min}})\gamma+2s_{\text{min}}}.$ (45)
This formulation is similar to the polarity objective in Eq.  2, which instead uses the denominator $|S_{0}|+|S_{1}|$. Like polarity, $\gamma$-polarity penalizes large or imbalanced clusters through a normalization term related to cluster sizes. However, the specific form of the denominator in Eq. 45 additionally enforces balance between $S_{1}$ and $S_{2}$.
In contrast, our objective promotes balanced clustering through an additive regularization term which supports an arbitrary number of clusters $k$. A key advantage of additive regularization is its simplicity: it effectively corresponds to shifting intra-cluster similarities by $-\beta$ (see Prop. 1). This straightforward modification enables us to establish a linear convergence rate (Thm. 3) and significantly improves computational efficiency (see Section 3.4). In comparison, obtaining similar theoretical guarantees and efficiency via local search on polarity-based objectives, including $\gamma$-polarity, is likely infeasible or substantially more complex.
In summary, our objective facilitates the derivation of a local search algorithm for the PCD problem that scales efficiently to large graphs.
Appendix G Experiments: More Details and Further Results
All experiments were conducted locally on a single machine with an Intel Core i9-10850k CPU and 64 GB of RAM.
G.1 Datasets
@@
Following [44], we consider the following widely studied real-world signed networks. WoW-EP8 (W8) [25] represents interactions among authors in the 8th EU Parliament legislature, where edge signs indicate collaboration or competition. Bitcoin (BTC) [31] is a trust-distrust network of users trading on the Bitcoin OTC platform. WikiVot (WikiV) [31] records positive and negative votes for Wikipedia admin elections. Referendum (REF) [30] captures tweets about the 2016 Italian constitutional referendum, with edge signs indicating whether users share the same stance. Slashdot (SD) [31] is a friend-foe network from the Slashdot Zoo feature. WikiCon (WikiC) [26] tracks positive and negative interactions between users editing English Wikipedia. Epinions (EP) [31] represents the trust-distrust relationships in the Epinions online social network. WikiPol (WikiP) [33] captures interactions among users editing Wikipedia pages on political topics.
G.2 Baselines
@@
For SCG, we use the public implementation from [40]. For KOCG, we use the public implementation from [32] with default hyperparameters: $\alpha=1/(k-1)$, $\beta=50$ (note that the purpose of this $\beta$ differs from the one used in our paper), and $\ell=5000$. For the spectral methods SPONGE and BNC, we use the public implementations from [37]. Following [44], for SPONGE, we evaluate both the unnormalized and symmetric normalized versions and report results for the best-performing method. For N2PC [21], we use default parameters for training their framework (based on GNN). We use the public implementation provided by the authors.
G.3 Description of Synthetic Datasets
We employ the modified signed stochastic block model (m-SSBM), which was specifically designed to generated synthetic graphs with planted ground-truth communities for PCD. The m-SSBM model is parameterized by four variables: (i) $n$, the total number of nodes; (ii) $k$, the number of non-neutral clusters; (iii) $\ell$, the size of each non-neutral cluster; and (iv) $\eta\in[0,1]$, which controls the edge probabilities. Edges within the same cluster are positive with probability $1-\eta$, and negative or absent with probability $\eta/2$. Conversely, edges between different clusters are negative with probability $1-\eta$, and positive or absent with probability $\eta/2$. All other edges are positive or negative with equal probability $\min(\eta,1/2)$.
Smaller values of $\eta$ correspond to denser non-neutral clusters and lower levels of noise. In other words, $\eta$ controls both sparsity of the graph and noise (i.e., flipping sign of edge weights).
G.4 Results on Synthetic Datasets with Imbalanced Clusters
Figure 2: F1-score and polarity of different methods on synthetic graphs generated using the m-SSBM model, as the size ratio parameter $\rho$ varies. A larger value of $\rho$ means the ground-truth non-neutral clusters are more imbalanced. The noise level is fixed to $\eta=0.4$.
We now evaluate the performance of different methods on synthetic data with imbalanced cluster sizes. We use the same m-SSBM model described in the main text (and previous subsection). To control the degree of imbalance among the $k$ planted groups, we follow the approach of [23] and introduce a group size ratio parameter $\rho\geq 1$. When $\rho=1$, each group has the same size $\ell$, resulting in a balanced partition. For $\rho>1$, the group sizes follow a geometric progression: the smallest group has size $s$, the next has size $s\cdot\rho^{1/(k-1)}$, and so on, such that the largest group is $s\cdot\rho$. This construction ensures that the ratio between the largest and smallest group sizes is exactly $\rho$, while the total number of non-neutral nodes remains approximately $k\ell$. Any remaining nodes (i.e., when $n>\sum_{i=1}^{k}|C_{i}|$) are assigned as neutral nodes, and their edges are sampled from a neutral distribution. In short, a larger value of $\rho$ means more imbalanced ground-truth non-neutral clusters.
The results are presented in Fig. 2. The noise level is fixed to $\eta=0.4$. We observe that our method remains robust as the cluster size imbalance increases, while the performance of baseline methods deteriorates more rapidly. This suggests that our approach does not rely on strict balance assumptions and is capable of effectively recovering imbalanced ground-truth clusterings.
G.5 Runtime Comparison of Local Search Improvements
Figure 3: Comparison of runtime for the three implementations of Alg. 2 (local search) introduced in Section 3.4 by varying the graph size $n$ and the number of non-neutral clusters $k$, using data generated from the m-SSBM model. The noise level is fixed at $\eta=0.4$. When varying $n$, we fix $k=4$; when varying $k$, we fix $n=5000$.
In Section 3.4, we introduced three different implementations of the local search procedure in Alg. 2: (i) the naive approach, referred to as LSNaive, which has a computational complexity of $O(k^{2}n^{2})$; (ii) an improved version that leverages the gradient, reducing the complexity to $O(Tkn)$, which we call LSImproved; and (iii) the final optimized version—used as LSPCD in all our experiments—presented in Alg. 3, which achieves a complexity of $O(kn^{2}+T(n+k))$.
Fig. 3 presents the runtime comparison as we vary the number of nodes $n$ and the number of non-neutral clusters $k$, using datasets generated from the m-SSBM model (Section G.3) with fixed noise level $\eta=0.4$. We fix $k=4$ when varying $n$, and $n=5000$ when varying $k$. LSNaive is included only for small values of $n$ due to its high computational cost, as shown in the leftmost plot. Across the remaining plots, the method in Algorithm 3 (LSPCD) demonstrates superior scalability with respect to both $n$ and $k$. While these are moderate-sized graphs, the performance gap would widen significantly for larger instances.
G.6 Aspects to Assess Solution Quality
@@
Below, we present 11 evaluation criteria used in our experiments on real-world data to shed light on how the clustering solutions generated by each method differ. In the context of PCD (where ground-truth solutions are not available), evaluating clustering quality is inherently subjective: each aspect listed below highlights a distinct facet of the solution. Generally, there is a trade-off among these aspects, and the objective is to achieve a good balance between them. Below, we define the aspects used in our analysis. We have defined these aspects such that a larger number is better (apart from runtime).
@@
  • IF: The imbalance factor introduced in the main text.
  • POL: Polarity, as defined in the main paper (Eq. 2).
@@
  • TIME (s): Runtime of the corresponding method in seconds.
G.7 Varying $\alpha$ and $\beta$
Fig. 4 shows the effect of varying $\beta$. Very small or large $\beta$ values lead to poorer polarity, as they produce clustering solutions with too many or too few non-neutral objects, respectively. In contrast, intermediate $\beta$ values consistently yields competitive polarity with the best performing methods, while being more balanced.
In Fig. 5, we illustrate the impact of varying $\alpha$ and $\beta$. According to Fig. 5(a), increasing $\alpha$ naturally balances intra-cluster cohesion and inter-cluster opposition. The size proportion, defined as the fraction of non-neutral objects in $V$, remains constant as $\alpha$ varies. Fig. 5(b) shows that increasing $\beta$ monotonically reduces the number of non-neutral objects, leading to denser clusters, as indicated by improved MAC and MAO scores. Notably, balance remains stable across different $\beta$ values, unlike the baseline SCG-MA.
Figure 4: Impact of $\beta$ on the WikiPol dataset with $k=6$.
(a) Epinions, $k=2$, $\beta=0.1$
(b) Wikipol, $k=6$, $\alpha=\frac{1}{k-1}$
Figure 5: Investigation of the impact of $\alpha$ and $\beta$.
G.8 Results on Real-World Datasets
Tables 4-11 present detailed analyses for all datasets. See Appendix G.6 for a description of the 11 aspects reported. Firstly, we observe that our method exhibits low standard deviation, indicating robustness to the initial random solution. It consistently finds high-polarity solutions while maintaining better balance than its main competitor, SCG. Unlike some baselines such as KOCG and SPONGE, our method does not enforce excessive balance, ensuring solutions remain both high in polarity and reasonably balanced.